OpenStudy (anonymous):

If the puppies’ parents have long hair, then the puppies will have long hair. What is the underlined portion of the statement called?

OpenStudy (anonymous):

what is it? I don't even

OpenStudy (anonymous):

Kena you're going to have to tell us which portion of the question is supposed to be underlined.

OpenStudy (anonymous):

lol soryy "then the puppies will have long hair"

OpenStudy (anonymous):

also, isn't this english? haha

OpenStudy (anonymous):

idk its a math quiz but this question is on it

OpenStudy (anonymous):

That portion of the statement is the consequent. It goes antecedent -> consequent.

OpenStudy (anonymous):

lol sounds like your teacher is messing with you.

OpenStudy (anonymous):

This is mathematical logic, not english.

OpenStudy (anonymous):

conclusion hypothesis counterexample assumption these are the answers listed???

OpenStudy (anonymous):

Conclusion, then.

OpenStudy (anonymous):

no, assumtion. Just because the parents have long hair does NOT mean the puppies will.

OpenStudy (anonymous):

bl4ckj4ck777, please don't answer questions that you don't know the answer to.

OpenStudy (anonymous):

thanks do you guys know the answer to this??? Find the counterexample to this statement: If ab = 3, then either a or b must be equal to 3.

OpenStudy (anonymous):

a=-1, b=-3

OpenStudy (anonymous):

excuse me, but it's true, there is no conclusion because nothing has happened yet. Now that I have the information to answer the question, I can. Genetics says that there are dominant and recessive genes, and just because there are two dominant genes bein portrayed does not mean that the resulting puppies will have the dominant trait so long as both parents ALSO have the recessive genes.

OpenStudy (anonymous):

This isn't a genetics question. It's a logic question. The word conclusion means something different in logic than it does in science at large. Again, please, if you don't know the subject that the question is taken from, allow someone who does to answer it.

OpenStudy (anonymous):

1/3 ·3 = 1 1/3·9 = 3 3·2 = 6 3·1 = 3 these are the answers available???

OpenStudy (anonymous):

1/3 x 9 = 3 works.

OpenStudy (anonymous):

what about this question?? Identify whether the following statement is inductive or deductive reasoning: You observe that whenever you divide the circumference of a circle by its diameter, you get a value close to . You conclude that the ratio of the circumference of a circle to its diameter is .

OpenStudy (anonymous):

I guess I just misunderstood the question then, I'm sorry that I didn't realize it was a specific genre of math/logic. I was answering the question to the best of my knowledge, not trying to throw a bs answer.

OpenStudy (anonymous):

kena, it's best to open a new question for each new question that you have. That question is inductive reasoning. Inductive reasoning is where you have multiple examples of a thing and you infer a general truth from it. Deductive reasoning is where you have a general truth, and you apply it to a situation to gain a specific knowledge.